# Anomaly-induced vanishing of brane partition functions

### v1

@@ 6.1 Trivial fibration
For $v_{3}=w_{2}w_{1}$, we again utilise the fact that Wu classes vanish at degrees above half the dimension of the manifold, so $V_{4}=\beta(v_{3})=0$ implies $v_{3}$ admits an integral lift. We would like to argue that, in this particular context, one can use $W_{3}$ as an effective integral lift of $v_{3}$. To see how, after dimensional reducing (157) over $T^{2}$ to a 4d action, we will find terms of the form $x_{1}\cup w_{2}\cup w_{1}$, where $x_{1}\in H^{1}(X^{4};\mathbb{Z}_{2})$ roughly corresponds either to $b_{1}]_{\mathbb{R}/\mathbb{Z}}$ or $c_{1}]_{\mathbb{R}/\mathbb{Z}}$ in (174). Recall that from the perspective of anomaly inflow, the anomaly theory of the D3-brane is defined on a 6-manifold $Z^{6}$ with $\partial Z^{6}=X^{4}\times S^{1}$. In this viewpoint, $x_{1}$ originates from a 3-cocycle $x_{3}\in H^{3}(Z^{6};\mathbb{Z}_{2})$. It follows from the Cartan formula that
@@
where the Adem relations are used in the second equality. The integral uplift of $\text{Sq}^{3}:H^{\ast}(-;\mathbb{Z}_{2})\to H^{\ast+3}(-;\mathbb{Z}_{2})$ is precisely the differential $d_{3}=\beta\circ\text{Sq}^{2}\circ(\text{mod}\ 2)$ on the third page of the Atiyah-Hirzebruch spectral sequence for (complex) K-theory. Hence, $\text{Sq}^{3}(x_{n-3})=0$ corresponds to the condition for $x_{n-3}$ to admit a K-theory lift, whereas $W_{3}=0$ (plus $w_{1}=0$) is the condition for a manifold to be orientable in K-theory (see, e.g. Diaconescu:2000wy ; Freed:2000ta ; Yonekura:2024bvh ).
@@ 6.2 Non-trivial fibration with constant local system
We shall now consider a non-trivial elliptic fibration $T^{2}\xrightarrow{\iota}X^{6}\xrightarrow{\pi}X^{4}$ which is not a product manifold. The base $X^{4}$ can be allowed to be not simply-connected, i.e. $\pi_{1}(X^{4})\neq 0$, but we demand that it acts trivially on the fiber $T^{2}$, such that the system of local coefficients is constant. We will treat the most general case in the next subsection. To obtain the basepoint anomaly of the D3-brane on $X^{4}$, starting from that of the M5-brane on $X^{6}$, we need to perform a fiber integration over $T^{2}$. Formally, the mathematical tool relating the cohomology groups between the total space, the base, and the fiber is the Leray-Serre spectral sequence 1b68bd4a-2ddc-338a-95d3-30d18474da26 . The primary input data for the spectral sequence are entries on the second page,
@@
where $T_{\pi}X=\text{ker}(d\pi)$ denotes the vertical tangent bundle, i.e. the bundle of vectors in $TX$ tangent to the fibers 2a7f5c42-423c-3a9f-b9ae-23d812c6429c , so the Stiefel-Whitney classes of the total space decompose as
@@ 6.3 Non-trivial fibration with non-trivial local system
where $\mathcal{F}]_{\mathbb{Z}^{2}_{\rho}}\in H_{0}(X^{4};\mathbb{Z}^{2}_{\rho})$ is taken to be a sum of the independent generators of $H_{0}(X^{4};\mathbb{Z}^{2}_{\rho})$, up to a preferred choice of normalisation. Note that the cap product is taken with respect to the ordinary (co)homology of the total space $X^{6}$, so it simply outputs an element of $\mathbb{Z}$ without twists.
@@
Taking inspiration from 10.1215/kjm/1250517912 ; Greenblatt2006Homology , we postulate a generalisation of the Stiefel-Whitney classes,
@@
To motivate our proposal, consider the case where $\rho$ acts on $\mathbb{Z}\oplus\mathbb{Z}$ simply by a sign flip on both copies. Such a local system factorises as $(\mathbb{Z}\oplus\mathbb{Z})_{\rho}\cong\widetilde{\mathbb{Z}}\oplus\widetilde{\mathbb{Z}}$, where $\widetilde{\mathbb{Z}}\coloneqq\mathbb{Z}_{w_{1}}$ is the orientation module corresponding to a non-trivial first Stiefel-Whitney class $w_{1}$. Meanwhile, $\rho$ acts trivially on $\mathbb{Z}_{2}$, i.e. $(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho}\cong\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, so we recover a doublet $\mathfrak{w}_{i}=(w_{i},w_{i})$ with $w_{i}\in H^{i}(B\text{O}(n);\mathbb{Z}_{2})$ being the standard $i$-th Stiefel-Whitney class. The cohomology ring $H^{\ast}(B\text{O}(n);\mathbb{Z}_{w_{1}})$ is indeed non-trivial 10.1215/kjm/1250517912 , such that $\mathcal{W}_{i+1}=\beta(\mathfrak{w}_{i})$ measures the obstruction to find twisted integral lifts of $\mathfrak{w}_{i}$.
@@
for any $\mathfrak{a}_{n-i}\in H^{n-i}(M^{n};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho})$, where $\mathfrak{Sq}^{i}:H^{\ast}(M^{n};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho})\to H^{\ast+i}(M^{n};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho})$ are regarded as cohomology operations with local coefficients 1c7ecebe-252f-3c02-8b70-bd46894bf791 satisfying axioms analogous to those of the ordinary version. Furthermore, assuming they are related to the twisted Stiefel-Whitney classes by the Wu formula
@@ 6.4 D3-brane on S-folds
As a concrete application, we would like to study the behavior of the D3-brane partition function in a class of non-trivial F-theory backgrounds known as S-folds. These are generalisations of orientifolds in Type IIB string theory Witten:1998xy . The latter are holographically dual to 4d $\mathcal{N}=4$ SCFTs, while the former are dual to 4d $\mathcal{N}=3$ SCFTs Garcia-Etxebarria:2015wns ; Aharony:2016kai . [^37] See also Apruzzi:2020pmv ; Giacomelli:2020gee for a construction of $\mathcal{N}=2$ S-folds.
@@
where $C^{i}\cong A$ and $t$ is a generator of $\mathbb{Z}_{k}$ Davis2001LectureNI ; MR1867354 ; Aharony:2016kai ; Heckman:2022muc ; Etheredge:2023ler . By construction, $1+t+\cdots+t^{k-1}=0$, and $d\circ d=(1-t)(1+t+\cdots+t^{k-1})=1-t^{k}=0$, i.e. the differential is nilpotent, as desired.
@@ 6.4.1 Discrete torsion and twisted Stiefel-Whitney classes
Apart from the integer $k$, an S-fold is also characterised by a choice of discrete torsion $\Theta=(\Theta_{\text{NS}},\Theta_{\text{RR}})\in H^{3}(S^{5}/\mathbb{Z}_{k};(\mathbb{Z}\oplus\mathbb{Z})_{\rho_{k}})$ Aharony:2016kai , which determines the cohomology class of $\mathcal{H}_{3}=(H_{3},G_{3})$ in the supergravity background. For example, when $k=2$, the gauge group of the 4d SCFT dual to the supergravity background is $\text{Sp}(M)$ or $\text{SO}(M)$ for some $M\in\mathbb{Z}^{+}$, depending on whether $\Theta_{\text{NS}}$ is non-trivial. When $\Theta_{\text{NS}}$ is trivial, the rank of $\text{SO}(M)$ depends on whether $\Theta_{\text{RR}}$ is non-trivial Witten:1998xy . Suppose we wrap a D3-brane on, say, $S^{1}\times S^{3}/\mathbb{Z}_{k}\subset\text{AdS}_{5}\times S^{5}/\mathbb{Z}_{k}$, our question is whether its partition function is vanishing or not in the presence of a non-trivial discrete torsion. Equivalently, we know from (209) that we need the 4-manifold $S^{1}\times S^{3}/\mathbb{Z}_{k}$ to have a non-trivial $\mathcal{W}_{3}$ that cancels $\mathcal{H}_{3}$. [^38] For the twisted fiber integration (200) to work, we require $E_{\infty}^{5,0}=0$ and and $H_{0}(X^{4};\mathbb{Z}^{2}_{\rho})$ to be non-trivial, as explained in Appendix B. The former is always satisfied because of degree reasons, while the latter indeed holds for our lens spaces, except for the trivial case of $k=6$.
@@
For $k=2$, we observe that $\rho_{2}$ corresponds to the identity matrix in $\text{SL}(2,\mathbb{Z}_{2})$, because a sign flip mod 2 is not meaningful. The local system is then trivial when considering $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ coefficients. In this case, $\mathfrak{w}_{2}=(w_{2},w_{2})\in H^{2}(S^{3}/\mathbb{Z}_{2};\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})$ is simply two identical copies of the second Stiefel-Whitney class. Provided that $S^{3}/\mathbb{Z}_{2}\cong\mathbb{RP}^{3}$ is spin, and so $\mathcal{W}_{3}=\beta(\mathfrak{w}_{2})=0$, we again conclude the D3-brane partition function must be vanishing when the discrete torsion $\Theta\in\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ is non-trivial, which matches with the analysis of the orientifold in Witten:1998xy .
@@ 6.4.2 Non-Abelian corrections
Since $H^{2}(S^{3}/\mathbb{Z}_{k};(\mathbb{Z}\oplus\mathbb{Z})_{\rho_{k}})=0$, any non-trivial element $\xi_{2}=(\zeta_{2}^{B},\zeta_{2}^{C})\in H^{2}(S^{3}/\mathbb{Z}_{k};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho_{k}})$ cannot be a mod $m$ reduction, which implies $\beta(\xi)\neq 0$ by exactness. The corresponding stack of D3-branes then has a non-vanishing partition function only if
@@
By the same token, for $k=4$, we have $H^{2}(S^{3}/\mathbb{Z}_{4};(\mathbb{Z}_{m}\oplus\mathbb{Z}_{m})_{\rho_{4}})=\mathbb{Z}_{\gcd(m,2)}$, so we can take any $N\in 2\mathbb{Z}$ and $m\in 2\mathbb{Z}$ a divisor of $N$ to match the discrete torsion $\Theta\in\mathbb{Z}_{2}$. This generalises the specific $N=4$ example studied in Aharony:2016kai to an infinite family of candidates. The trivial case of $k=6$ is uninteresting as usual. All in all, from the perspective of obstruction theory, we have constructed non-Abelian D3-brane configurations on S-folds which have non-vanishing partition functions in the presence of discrete torsion.
@@ 6.5 Class $\mathcal{S}$ theories
We briefly outline a generalisation that is applicable to the compactification of the M5-brane worldvolume theory, viewed as a 6d $(2,0)$ SCFT, over a generic Riemann surface $\Sigma^{2}$, which results in what are referred to as 4d Class $\mathcal{S}$ theories Gaiotto:2009hg (see also Lawrie:2018jut ). These constructions typically preserve $\mathcal{N}=2$ supersymmetry, and often admit no known Lagrangian descriptions Moore:2012yp ; Tachikawa:2013kta ; Akhond:2021xio . For the sake of illustration, let us assume below that the Riemann surface is compact, orientable, and has no punctures, so $\Sigma^{2}$ is characterised only by its genus $g$.
@@ Acknowledgements.
We thank Victor Carmona, Dan Freed, Andrea Grigoletto, Patrick Jefferson, Nitu Kitchloo, Ho Tat Lam, Daniel Roggenkamp, Raffaele Savelli, Sebastian Schulz, and Thomas Waddleton for useful discussion and correspondence. EL is grateful to the Department of Mathematical Sciences at Durham University for hospitality during a visit contributing to this work. The work of IGE is supported by the “Global Categorical Symmetries” Simons Collaboration grant (award number 888990) and the STFC ST/X000591/1 consolidated grant.

### v3

@@ 6.1 Trivial fibration
For $v_{3}=w_{2}w_{1}$, we again utilise the fact that Wu classes vanish at degrees above half the dimension of the manifold, so $V_{4}=\beta(v_{3})=0$ implies $v_{3}$ admits an integral lift. We would like to argue that, in this particular context, one can use $W_{3}$ as an effective integral lift of $v_{3}$. To see how, after dimensionally reducing (157) over $T^{2}$ to a 4d action, we will find terms of the form $x_{1}\cup w_{2}\cup w_{1}$, where $x_{1}\in H^{1}(X^{4};\mathbb{Z}_{2})$ roughly corresponds either to $b_{1}]_{\mathbb{R}/\mathbb{Z}}$ or $c_{1}]_{\mathbb{R}/\mathbb{Z}}$ in (174). Recall that from the perspective of anomaly inflow, the anomaly theory of the D3-brane is defined on a 6-manifold $Z^{6}$ with $\partial Z^{6}=X^{4}\times S^{1}$. In this viewpoint, $x_{1}$ originates from a 3-cocycle $x_{3}\in H^{3}(Z^{6};\mathbb{Z}_{2})$. It follows from the Cartan formula that
@@
where the Adem relations are used in the second equality. The integral uplift of $\text{Sq}^{3}:H^{\ast}(-;\mathbb{Z}_{2})\to H^{\ast+3}(-;\mathbb{Z}_{2})$ is precisely the differential $d_{3}=\beta\circ\text{Sq}^{2}\circ(\text{mod}\ 2)$ on the third page of the Atiyah-Hirzebruch spectral sequence for (complex) K-theory. Hence, $\text{Sq}^{3}(x_{n-3})=0$ corresponds to the condition for $x_{n-3}$ to admit a K-theory lift, whereas $W_{3}=0$ (plus $w_{1}=0$) is the condition for a manifold to be orientable in K-theory (see, e.g. Diaconescu et al. (2003a); Freed (2000); Yonekura (2024)).
@@ 6.2 Non-trivial fibration with constant local system
We shall now consider a non-trivial elliptic fibration $T^{2}\xrightarrow{\iota}X^{6}\xrightarrow{\pi}X^{4}$ which is not a product manifold. The base $X^{4}$ can be allowed to be not simply-connected, i.e. $\pi_{1}(X^{4})\neq 0$, but we demand that it acts trivially on the cohomology of the fiber $T^{2}$, such that the system of local coefficients appearing in the spectral sequence below is constant. We will treat the most general case in the next subsection. To obtain the basepoint anomaly of the D3-brane on $X^{4}$, starting from that of the M5-brane on $X^{6}$, we need to perform a fiber integration over $T^{2}$. Formally, the mathematical tool relating the cohomology groups between the total space, the base, and the fiber is the Leray-Serre spectral sequence Serre (1951). The primary input data for the spectral sequence are entries on the second page,
@@
where $T_{\pi}X=\text{ker}(d\pi)$ denotes the vertical tangent bundle, i.e. the bundle of vectors in $TX$ tangent to the fibers Szczarba (1964), so the Stiefel-Whitney classes of the total space decompose as
@@ 6.3 Non-trivial fibration with non-trivial local system
where $\mathcal{F}]_{\mathbb{Z}^{2}_{\rho}}\in H_{0}(X^{4};\mathbb{Z}^{2}_{\rho})$ is taken to be a sum of the independent generators of $H_{0}(X^{4};\mathbb{Z}^{2}_{\rho})$, up to a preferred choice of normalisation. [^36] A more precise definition of $\mathcal{F}]_{\mathbb{Z}^{2}_{\rho}}$ is provided in Appendix B. It is indeed not always guaranteed that such a class can be constructed and satisfies the assumptions therein, but we have checked that it is possible to do so in the examples considered in this paper. Note that the cap product is taken with respect to the ordinary (co)homology of the total space $X^{6}$, so it simply outputs an element of $\mathbb{Z}$ without twists.
@@
Taking inspiration from Čadek (1999); Greenblatt (2006), we postulate a generalisation of the Stiefel-Whitney classes,
@@
To motivate our proposal, consider the case where $\rho$ acts on $\mathbb{Z}\oplus\mathbb{Z}$ simply by a sign flip on both copies. Such a local system factorises as $(\mathbb{Z}\oplus\mathbb{Z})_{\rho}\cong\widetilde{\mathbb{Z}}\oplus\widetilde{\mathbb{Z}}$, where $\widetilde{\mathbb{Z}}\coloneqq\mathbb{Z}_{w_{1}}$ is the orientation module corresponding to a non-trivial first Stiefel-Whitney class $w_{1}$. Meanwhile, $\rho$ acts trivially on $\mathbb{Z}_{2}$, i.e. $(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho}\cong\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, so we recover a doublet $\mathfrak{w}_{i}=(w_{i},w_{i})$ with $w_{i}\in H^{i}(B\text{O}(n);\mathbb{Z}_{2})$ being the standard $i$-th Stiefel-Whitney class. The cohomology ring $H^{\ast}(B\text{O}(n);\mathbb{Z}_{w_{1}})$ is indeed non-trivial Čadek (1999), such that $\mathcal{W}_{i+1}=\beta(\mathfrak{w}_{i})$ measures the obstruction to find twisted integral lifts of $\mathfrak{w}_{i}$.
@@
for any $\mathfrak{a}_{n-i}\in H^{n-i}(M^{n};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho})$, where $\mathfrak{Sq}^{i}:H^{\ast}(M^{n};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho})\to H^{\ast+i}(M^{n};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho})$ are regarded as cohomology operations with local coefficients Gitler (1963) satisfying axioms analogous to those of the ordinary version. Furthermore, assuming they are related to the twisted Stiefel-Whitney classes by the Wu formula
@@ 6.4 D3-brane on S-folds
As a concrete application, we would like to study the behavior of the D3-brane partition function in a class of non-trivial F-theory backgrounds known as S-folds. These are generalisations of orientifolds in Type IIB string theory Witten (1998). The latter are holographically dual to 4d $\mathcal{N}=4$ SCFTs, while the former are dual to 4d $\mathcal{N}=3$ SCFTs García-Etxebarria and Regalado (2016); Aharony et al. (2016). [^38] See also Apruzzi et al. (2020); Giacomelli et al. (2021) for a construction of $\mathcal{N}=2$ S-folds.
@@
where $C^{i}\cong A$ and $t$ is a generator of $\mathbb{Z}_{k}$ Davis and Kirk (2001); Hatcher (2002); Aharony et al. (2016); Heckman et al. (2023); Etheredge et al. (2023). By construction, $1+t+\cdots+t^{k-1}=0$, and $d\circ d=(1-t)(1+t+\cdots+t^{k-1})=1-t^{k}=0$, i.e. the differential is nilpotent, as desired.
@@ 6.4.1 Discrete torsion and twisted Stiefel-Whitney classes
Apart from the integer $k$, an S-fold is also characterised by a choice of discrete torsion $\Theta=(\Theta_{\text{NS}},\Theta_{\text{RR}})\in H^{3}(S^{5}/\mathbb{Z}_{k};(\mathbb{Z}\oplus\mathbb{Z})_{\rho_{k}})$ Aharony et al. (2016), which determines the cohomology class of $\mathcal{H}_{3}=(H_{3},G_{3})$ in the supergravity background. For example, when $k=2$, the gauge group of the 4d SCFT dual to the supergravity background is $\text{Sp}(M)$ or $\text{SO}(M)$ for some $M\in\mathbb{Z}^{+}$, depending on whether $\Theta_{\text{NS}}$ is non-trivial. When $\Theta_{\text{NS}}$ is trivial, the rank of $\text{SO}(M)$ depends on whether $\Theta_{\text{RR}}$ is non-trivial Witten (1998). Suppose we wrap a D3-brane on, say, $S^{1}\times S^{3}/\mathbb{Z}_{k}\subset\text{AdS}_{5}\times S^{5}/\mathbb{Z}_{k}$, our question is whether its partition function is vanishing or not in the presence of a non-trivial discrete torsion. Equivalently, we know from (209) that we need the 4-manifold $S^{1}\times S^{3}/\mathbb{Z}_{k}$ to have a non-trivial $\mathcal{W}_{3}$ that cancels $\mathcal{H}_{3}$. [^39] For the twisted fiber integration (200) to work, we require $E_{\infty}^{5,0}=0$ and and $H_{0}(X^{4};\mathbb{Z}^{2}_{\rho})$ to be non-trivial, as explained in Appendix B. The former is always satisfied because of degree reasons, while the latter indeed holds for our lens spaces, except for the trivial case of $k=6$.
@@
For $k=2$, we observe that $\rho_{2}$ corresponds to the identity matrix in $\text{SL}(2,\mathbb{Z}_{2})$, because a sign flip mod 2 is not meaningful. The local system is then trivial when considering $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ coefficients. In this case, $\mathfrak{w}_{2}=(w_{2},w_{2})\in H^{2}(S^{3}/\mathbb{Z}_{2};\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})$ is simply two identical copies of the second Stiefel-Whitney class. Provided that $S^{3}/\mathbb{Z}_{2}\cong\mathbb{RP}^{3}$ is spin, and so $\mathcal{W}_{3}=\beta(\mathfrak{w}_{2})=0$, we again conclude the D3-brane partition function must be vanishing when the discrete torsion $\Theta\in\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ is non-trivial, which matches with the analysis of the orientifold in Witten (1998).
@@ 6.4.2 Non-Abelian corrections
Since $H^{2}(S^{3}/\mathbb{Z}_{k};(\mathbb{Z}\oplus\mathbb{Z})_{\rho_{k}})=0$, any non-trivial element $\xi_{2}=(\zeta_{2}^{B},\zeta_{2}^{C})\in H^{2}(S^{3}/\mathbb{Z}_{k};(\mathbb{Z}_{2}\oplus\mathbb{Z}_{2})_{\rho_{k}})$ cannot be a mod $m$ reduction, which implies $\beta(\xi_{2})\neq 0$ by exactness. The corresponding stack of D3-branes then has a non-vanishing partition function only if
@@
By the same token, for $k=4$, we have $H^{2}(S^{3}/\mathbb{Z}_{4};(\mathbb{Z}_{m}\oplus\mathbb{Z}_{m})_{\rho_{4}})=\mathbb{Z}_{\gcd(m,2)}$, so we can take any $N\in 2\mathbb{Z}$ and $m\in 2\mathbb{Z}$ a divisor of $N$ to match the discrete torsion $\Theta\in\mathbb{Z}_{2}$. This generalises the specific $N=4$ example studied in Aharony et al. (2016) to an infinite family of candidates. The trivial case of $k=6$ is uninteresting as usual. All in all, from the perspective of obstruction theory, we have constructed non-Abelian D3-brane configurations on S-folds which have non-vanishing partition functions in the presence of discrete torsion.
@@ 6.5 Class $\mathcal{S}$ theories
We briefly outline a generalisation that is applicable to the compactification of the M5-brane worldvolume theory, viewed as a 6d $(2,0)$ SCFT, over a generic Riemann surface $\Sigma^{2}$, which results in what are referred to as 4d Class $\mathcal{S}$ theories Gaiotto et al. (2013) (see also Lawrie et al. (2018)). These constructions typically preserve $\mathcal{N}=2$ supersymmetry, and often admit no known Lagrangian descriptions Moore (2012); Tachikawa (2013); Akhond et al. (2022). For the sake of illustration, let us assume below that the Riemann surface is compact, orientable, and has no punctures, so $\Sigma^{2}$ is characterised only by its genus $g$.
@@ Acknowledgements.
We thank T. Daniel Brennan, Victor Carmona, Dan Freed, Andrea Grigoletto, Patrick Jefferson, Nitu Kitchloo, Ho Tat Lam, Daniel Roggenkamp, Raffaele Savelli, Sebastian Schulz, and Thomas Waddleton for useful discussion and correspondence. EL is grateful to the Department of Mathematical Sciences at Durham University for hospitality during a visit contributing to this work. The work of IGE is supported by the “Global Categorical Symmetries” Simons Collaboration grant (award number 888990) and the STFC ST/X000591/1 consolidated grant.